1.b is defined by bn = n(−1)n, n ≥ 1

(a) Find Σ when 4 is on top, i=1 at the bottom and bi at the right hand side

(b) Is b increasing, decreasing, non-increasing or non-decreasing?


1
Expert's answer
2021-11-01T19:56:34-0400

(a)


"\displaystyle\sum_{i=1}^{4}v_i=v_1+v_2+v_3+v_4"

"=1(-1)^1+2(-1)^2+3(-1)^3+4(-1)^4"

"=-1+2-3+4=2"

(b)


"b_{2k-1}=(-1)^{2k-1}(2k-1)=-(2k-1)<0, k\geq1"

"b_{2k}=(-1)^{2k}(2k)=2k>0, k\geq1"

The sequence "\{b_n\}" is alternating.

The alternating sequence "\{b_n\}" is neither increasing nor decreasing .